OpenStudy (anonymous):

Solve S = 4v2 for v. Solve M = 2x + 3y for y.

OpenStudy (anonymous):

ok for S you can just isolate the v would you would divide both by 4 to make S/4 = v^2. then square root to get rid of the square of the v = \[\sqrt{S/4}\]

OpenStudy (anonymous):

to solve for y you kind of do the same thing and isolate the y. do this by minus the 2x from both sides to give M-2x = 3y then divide 3 by both sides. with this you get \[M/3 - 2x/3 = y\]

OpenStudy (anonymous):

ok thank you . so does that mean that what you wrote is the solution?

OpenStudy (anonymous):

yes. to both of them. and how to solve them in case you run up on another one you can do it for yoruself

OpenStudy (anonymous):

ok :) thank you